Please help me solve this Basic Counting Principle problem:
Consider the following.
(a) In how many different ways can a player select a club and a diamond from a deck of 52 bridge cards?
ways
(b) In how many different ways can a player select one card of each suit from a bridge deck?
ways

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

(a) In how many different ways can a player select a club and a diamond from a deck of 52 bridge cards?
------
A club is 1 of 4 = 1/4
Then the diamond is 13/51
The product is 13/204
=======================
(b) In how many different ways can a player select one card of each suit from a bridge deck?
The 1st is 1/4
Then 13/51
Then 13/50
Then 13/49
Multiply those.

Answer by greenestamps(13209)



Neither problem is clearly defined. I will assume....

(1) without replacement
(2) order does not matter

(a) a club and a diamond

On the first draw, there are 52 cards to choose from, of which 26 are okay (13 clubs and 13 diamonds). P = 26/52 = 1/2.

On the second draw, there are 51 cards remaining, of which 13 are okay (the 13 of the suit that was not drawn on the first draw). P = 13/51.

ANSWER: (1/2)(13/51) = 13/102

(b) one card of each suit

1st draw: 52 cards; can choose any. P = 52/52 = 1
2nd draw: 51 cards remaining; can choose any of the 39 in the 3 remaining suits. P = 39/51 = 13/17
3rd draw: 50 cards remaining; can choose any of the 26 cards in the 2 remaining suits. P = 26/50 = 13/25
4th draw: 49 cards remaining; must choose one of the 13 cards in the last remaining suit. P = 13/49

ANSWER: (1)(13/17)(13/25)(13/49)

Use a calculator and express the result in the desired/required form.